# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ V.6 Inference prospects for asymmetric-mass binaries with LISA
We remind the reader that all parameter inference was conducted using second-generation TDI variables as implemented by the fastLISAresponse software katz2022assessing , assuming constant- and equal-armlength orbits; we sampled all posterior distributions with Markov chain Monte-Carlo (MCMC) using the eryn package Karnesis:2023ras ; michael_katz_2023_7705496 ; Foreman-Mackey:2013 and default settings.
@@
$1$ $9.181$ $9.181$ 0.023 0.785 3.802 $1.049$ (28)
$2$ 16.506 16.505 0.040 0.295 5.326 $3.893$ (29)
$3$ 2.459 2.455 2.600 2.545 0.776 $17.006$ (32)
$4$ 2.650 2.650 5.234 69.845 0.870 $1.079$ (30)
$5$ 9.200 9.200 11.830 0.154 5.847 $1.540$ (31)
Table 2: Relative precisions obtained in the parameter estimation of the sources listed in Table 1 (where $\delta x$ denotes the relative precision in parameter $x$, defined as the $1\,\sigma$ width of the marginal posterior on $x$ normalised by the median value). The (s) superscripts indicates that masses have been converted in the source frame. For each sample, we recover the redshift from the luminosity distance $d_{\mathrm{L}}$ assuming the PLANCK18 cosmology. We also report $\Delta\Omega^{\mathrm{sky}}$ (Eq. 16), which is $99\%$ of the sky-localisation area of each source in square degrees. In the last column, we provide the reference for the corresponding posterior distributions in Appendix F. Notably, $\delta m_{1}^{\mathrm{(s)}}$ and $\delta m_{2}^{\mathrm{(s)}}$ are a factor of $\sim 3$ orders of magnitude larger than the equivalent quantities computed in the detector frame, $\delta m_{i}^{\mathrm{(d)}}$. This is a consequence of the mass relation $m_{i}^{(d)}=(1+z_{\mathrm{r}})\,m_{i}^{(s)}$, which introduces the (much larger) uncertainty in $d_{\mathrm{L}}$ due to its relationship with $z_{\mathrm{r}}$ via our assumed cosmology.
@@ V.6.1 EMRIs
We investigate three EMRIs with mass ratios $\epsilon\in\{10^{-4},10^{-5},10^{-6}\}$ and parameters given by the fifth, first and second row of Table 1 respectively. As confirmed by our horizon redshift study in Section V.4, these sources are indeed observable and each are unique in their own way. Source 1 has high spin and initial eccentricity $(a,e_{0})=(0.998,0.730)$, and represents a fairly typical EMRI, assuming that many MBHs in nature are spinning rapidly Babak:2017tow ; Barausse:2012fy . The more extreme mass-ratio of Source 2 causes its inspiral trajectory to evolve on a time-scale much slower than that of Source 1. This is reflected by its two-year inspiral starting deep in the strong field of the primary ($p_{0}\sim 2.12$) for $(a,e_{0})=(0.998,0.425)$, terminating $\sim 0.15$ away from the separatrix. Notice that the final eccentricity $e_{f}\sim 0.261$; the orbit is still quite eccentric close to plunge, and therefore lies in the upper tail of the eccentricity-at-plunge population for EMRIs formed via capture in Ref. Babak:2017tow . The final EMRI we consider, (Source 5) is a retrograde inspiral that (due to its mass ratio $\epsilon=10^{-4}$) begins in the weak-field regime ($p_{0}=26.19$) with high $e_{0}=0.8$ and moderate spin $a=-0.5$, terminating near the separatrix with non-trivial eccentricity $e_{f}=0.195$. Based on a prior examination of EMRI astrophysics in Ref. Babak:2017tow , we believe these are characteristic asymmetric-mass binaries that are readily observable by LISA to plausible luminosity distances (and therefore redshifts).
@@
Measurement precisions on the parameters of these sources are shown in the corresponding rows of Table 2. For Source 1, we see that we can constrain the intrinsic parameters to usual levels ($\Delta\theta/\theta\sim 10^{-6}$) for a source with an SNR of $30$. From the marginal posteriors (shown in Fig. 28), we find that setting $\kappa=10^{-2}$ (retaining 235 modes) yields identical inference results to $\kappa=10^{-5}$ (1228 modes), with any systematic biases present being too small to identify. Similar results are obtained for Source 2 in terms of measurement precision (accounting for the lower SNR of this source), except for the improved determination of $e_{\mathrm{f}}$ which is expected due to the significant final eccentricity of this source. Correlations between parameters are more pronounced for this source than the other two, which is likely a feature of the mass ratio of this source leading to slower orbital parameter evolution. Similarly to Source 1, there is an absence of any systematic biases in Fig. 29 between the posteriors for $\kappa=10^{-5}$ (809 modes) and $\kappa=10^{-2}$ (186 modes). For Source 3, the spin constraint are wider than for the other two sources, which is expected as the effects of spin are less pronounced for this source (as the inspiral spends a large fraction of the two-year duration in the weaker field). The same can not be said for eccentricity effects (which are still significant in the weak field), which is reflected in the recovery of $e_{\mathrm{f}}$. Once again, no systematic biases in Fig. 31 are observable between posteriors for $\kappa=10^{-5}$ (928 modes) and $\kappa=10^{-2}$ (208 modes).
@@ V.6.2 IMRIs
As demonstrated in our horizon redshift study (Fig. 17), IMRIs will be observable by LISA with high SNRs to high redshifts. To our knowledge, we have developed the first waveform model capable of modelling highly-eccentric IMRIs into rapidly-spinning MBHs (at the adiabatic order) that is sufficiently rapid for use in Bayesian inference studies. To demonstrate these capabilities and perform an initial investigation of IMRI posterior structure, we examine two IMRIs with mass ratios $\epsilon=10^{-2}$ and parameters given by row 3 and 4 in Table 1. Due to the large mass ratio, we see that the orbital parameters evolve far more rapidly than for the EMRI sources we examined previously. Indeed, despite their high initial eccentricities $e_{0}=0.85$, the binaries have circularised significantly prior to plunge. Notably, the SNRs $\rho=\{500,200\}$ of these sources are also significantly larger, mainly due to the waveform amplitudes being proportional to $\mu\sim m_{2}$ (Eq. 1). However, having such high SNRs may lead to more significant systematic biases when waveform models with incomplete mode content are used in parameter inference, as we explore below.
@@
For both Source 3 and Source 4, waveforms with $\kappa=10^{-5}$ (949 and 626 modes respectively) are sufficiently complete for performing parameter estimation with no observable biases, as shown in Figs. 32 and 30. However, when attempting to infer the parameters with $\kappa=10^{-2}$ (not shown in Figures; 223 and 156 modes respectively) we observed statistically significant biases across all the IMRI parameters. Nearly all parameters were recovered with biases of $\gtrsim 3\,\sigma$ away from the truth (with $\sigma$ computed as the standard deviation of the posterior samples) that were readily identifiable compared to the scale of the posterior. Even for $\kappa=10^{-3}$ (412 and 291 modes respectively), we observe unacceptable biases in the luminosity distance and sky location on the order of $\sim 4\sigma$ and $\sim 7\sigma$ biases away from the true parameters respectively. Biases of this scale would propagate to population / cosmological analyses and may contaminate the results of these studies, especially when considering that luminosity distance biases will affect source-frame mass estimates as well. This behaviour is to be expected considering the higher SNRs of these sources — from Eq. 15, we can expect mismatches of $10^{-3}$ (which $\kappa=10^{-3}$ roughly leads to) to be insufficient for the analysis of signals with SNRs of hundreds. The fact that extrinsic parameters are most affected by the omission of impactful waveform modes is also expected, as this has been previously identified in the analysis of comparable-mass case binaries observed with LISA Yi:2025pxe .
@@
We report measurement precisions of parameters of astrophysical interest for these sources in Table 2. Despite the higher SNRs of these signals, their spins are recovered with larger uncertainties than for the prograde EMRI sources we examined. This can be understood in terms of the rapid evolution of the IMRI trajectories, which complete relatively few cycles in the strong-field regime where the effects of spin are most important. Eccentricity-at-plunge is also recovered more poorly; this has similar justification, in addition to the IMRI trajectories beginning at larger orbital separations and therefore entering the strong-field regime with lower eccentricities than the EMRI sources. Luminosity distances are recovered precisely (to better than $1\%$ precision) in both cases, which is to be expected as amplitude measurements depend strongly on the SNR. This has the interesting consequence that despite the recovery of detector-frame masses for the IMRI sources being less precise than for the EMRI sources, they still provide more precise recovery of source-frame masses due to the improvement in the luminosity distance measurement. Last, we note that the $99\%$ sky-localisation area for Source 3 ($\sim 17\,\mathrm{deg}^{2}$) is more than an order of magnitude larger than that of Source 4. This occurs because during early inspiral, the bulk of the mode spectrum for this source sits mainly at lower frequencies ($\lesssim 10^{-4}\,\mathrm{Hz}$) that are poorly measured by LISA. As the inspiral evolves during the observation, the mode spectrum shifts to higher frequencies (“entering” the lower end of the LISA band) and is measured well. While this still enables many parameters to be recovered precisely, the sky position of the source is determined by long-timescale variations in the LISA response function; as the waveform of this source only has significant SNR for part of the observation (such that less of these variations are measurable), this in turn leads to a poorer localisation of the source on the sky.
@@ VI.1 Summary of this work
By comparing with both the AAK and a quadrupolar variant of our waveform model, we demonstrated the importance of relativistic amplitudes and higher-order modes when peerforming SNR computations and parameter estimation. We found that SNRs computed with the AAK are more inaccurate for stronger-field high-mass systems, where the missing higher-order modes are most important and the semi-relativistic amplitudes of the AAK are least accurate. For all component masses considered, the AAK overestimated the SNR for large spins (by up to $60\%$ in the high-mass case), and exhibited larger SNR errors for larger eccentricities. This has ramifications for quantitative studies of the EMRI detection rate in the literature (such as Ref. Babak:2017tow ), which should be revisited with relativistic waveform models in order to obtain better rate estimates. Exploring inference biases under the linear signal approximation, we identify biases of $\sim 1\,\sigma$ with the AAK and $\sim 0.1\,\sigma$ with the quadrupolar relativistic model. While these biases may be identifiable in the parameter estimation of EMRIs, they are sufficiently small that these approximate models may be readily applicable in the search for EMRI signals in detector data (where any corresponding reductions in computational cost would be highly beneficial). An exploration of this possibility is left to future work.
To quantify and explore LISA’s sensitivity to GWs from asymmetric-mass binaries, we computed sky-averaged horizon (maximum) redshifts at which LISA can detect sources with varying component masses, spins and eccentricities. We found that LISA can detect EMRIs of mass ratio $10^{-6}$ ($10^{-4}$) at redshifts of up to 1 (5), and IMRIs of mass ratio $10^{-3}$ ($10^{-2}$) at redshifts of up to 10 (15), with only $\sim 10\%$ variations with respect to eccentricity. In the case of EMRIs with mass ratio $10^{-5}$, horizon redshifts varied from $0.5$ to $3$ as spin increased from 00 to $0.999$. For all configurations, we found that LISA was most sensitive to binaries with source-frame primary masses $m_{1}^{\mathrm{(s)}}\sim 10^{6}\,M_{\odot}$. Our horizon redshift results highlight that the detection and characterisation of asymmetric-mass binaries will enable LISA to effectively probe the evolution of MBH, intermediate-mass black hole and stellar-mass black hole populations over cosmic time, in line with LISA’s science objectives. The waveform model we have developed (as a component of the few framework) is unique in the respect that it is able to accurately reproduce the physical characteristics of this wide range of GW sources, providing the tools necessary for the exploration of LISA science with IMRI and EMRI observations.
@@ VI.2 Prospects for future work
Extension to generic orbits. The few framework is designed with generic, eccentric and precessing orbits in mind but some technical limitations must be overcome before such a model can be built. Of particular note is the increased difficulty in interpolating data for these inspirals; the parameter-space dimensionality increases to four which greatly expands the (already large) memory requirements for mode amplitude interpolation. Given that polar ($k\neq 0$) mode amplitudes must also be computed (increasing the total number of modes by more than an order of magnitude Fujita:2025pc ), extrapolating from the domain of our model leads to mode amplitude grids of $\sim 500\,\mathrm{GB}$ in size or more, which is several times too large to fit in the memory of existing GPU hardware. The overall cost of computing the underlying data for flux and amplitude interpolation also grows significantly as the higher dimensional parameter space necessitates sampling the space at more points. Furthermore, at each point in the parameter space the perturbation theory computation is more expensive as one must integrate over both an orbit’s radial and polar motions, and one must sum over both radial and polar frequency harmonics. This increases the computational cost of each mode which contributes to the flux by a factor of about 2020202030303030, and increases the number of modes which contribute to the flux by a similar factor. The problem remains highly parallelizable, so will not be difficult to exploit efficient codes which can compute these data Drasco:2005kz ; Fujita:2009us ; vandeMeent:2017bcc ; Chen:2023lsa ; Kerachian:2023oiw . However, the sheer quantity of data needed means that care is needed in designing the grid spacing across the parameter space before committing significant computational resources.
Inspiral-merger-ringdown models. For typical EMRIs, very little SNR is accumulated during the merger and ringdown portion of the waveform. With recent work showing that GSF results can also be used to model IMRIs LeTiec:2011bk ; vandeMeent:2020xgc ; Wardell:2021fyy ; Warburton:2021kwk , it becomes more important to include the merger and ringdown apte2019exciting ; Becker:2024xdi ; Lhost:2024jmw ; Leather:2025nhu . Previous approaches to computing these in perturbation theory involved computing the waveform using time-domain codes Taracchini:2014zpa ; Lim:2019xrb ; Strusberg:2025qfv and thus the results were not suitable for the few framework. To overcome this, recent work has cast the transition from inspiral to plunge and the plunging portions of the waveform within the multiscale framework Kuchler:2024esj ; Kuchler:2025hwx . Within this approach the online computation of the transition and the plunge involves solving new trajectory ODEs and the waveform is computed from a new set of amplitudes. Structurally, this is exactly the same as the present online inspiral computation and, as such, the results of these calculations can be readily incorporated into few.
@@
Post-adiabatic accuracy: including the second-order GSF. Current second-order GSF calculations are carried out via a multiscale framework Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb . Within this approach, the trajectory through the parameter space is computed using an extended version of Eq. 2 that takes the form
@@
The post-adiabatic forcing functions $\hat{f}^{(1)}_{\alpha}$ are very computationally expensive to compute, but they can be pre-computed and interpolated in the same manner as the adiabatic forcing functions. Second-order GSF calculations also compute the second-order waveform amplitudes. It is expected that these are not necessary to include in EMRI waveforms Burke:2023lno , but they may be required when modelling low-$q$ IMRIs to avoid biases in parameter estimation. In that case, the second-order amplitudes will be interpolated along the sparse trajectory during online waveform generation alongside the adiabatic amplitudes. This means that online waveform generation of post-adiabatic waveforms will be almost as fast as adiabatic models. At present, the $\hat{f}^{(1)}_{\alpha}$ are only known for quasi-circular inspirals into a non-rotating black hole Wardell:2021fyy , and these results are ready to be incorporated into few (as was initially investigated in Ref. Burke:2023lno ). It is an active area of research to compute $\hat{f}^{(1)}_{\alpha}$ for more complex orbital configurations Spiers:2023cip ; Leather:2023dzj ; Osburn:2022bby ; Wei:2025lva .
Post-adiabatic accuracy: including the spin on the secondary. The effect of the spin on the secondary enters the waveform phase at post-adiabatic order. In terms of waveform generation, a spinning secondary introduces another forcing term to Eq. 17 and the orbital frequencies are also modified. Both of these changes can readily be incorporated into few. At present, codes exist that can compute the necessary forcing terms for circular Harms:2015ixa ; Akcay:2019bvk ; Piovano:2020zin ; Mathews:2021rod ; Albertini:2024rrs , eccentric Skoupy:2021asz and generic orbits Skoupy:2023lih ; Drummond:2023wqc ; Piovano:2024yks .
@@
With the exception of GW memory and generic orbits, work is already underway on all of these extensions (and we note that few currently can generate generic inspirals using a $5$PN model Isoyama:2021jjd ). In addition to the above, few can also be extended to include environmental and beyond-GR physics. So long as any additional physics can be modelled as a forcing function added to Eq. 2 or Eq. 17 that do not depend on the orbital phase (or if such dependence can be transformed away VanDeMeent:2018cgn ; Lynch:2021ogr ; Lynch:2024hco ; Lynch:2024ohd ) it can readily be incorporated into our framework. As the few trajectory module now exists entirely in Python, if software for computing these modifications has already been developed it can be used directly to rapidly produce a new inspiral (and hence, waveform) model suitable for subsequent data analysis investigations. See, e.g., Refs. Copparoni:2025jhq and Speri:2024qak for first efforts in this direction.

### v2

@@ V.6 Inference prospects for asymmetric-mass binaries with LISA
We remind the reader that all parameter inference was conducted using second-generation time-delay interferometry (TDI) variables as implemented by the fastLISAresponse software katz2022assessing , assuming constant- and equal-armlength orbits; we sampled all posterior distributions with Markov chain Monte-Carlo (MCMC) using the eryn package Karnesis:2023ras ; michael_katz_2023_7705496 ; Foreman-Mackey:2013 and default settings.
@@
11 $9.181$ $9.181$ 0.023 0.785 3.802 $1.049$ (28)
22 16.506 16.505 0.040 0.295 5.326 $3.893$ (29)
33 2.459 2.455 2.600 2.545 0.776 $17.006$ (32)
44 2.650 2.650 5.234 69.845 0.870 $1.079$ (30)
55 9.200 9.200 11.830 0.154 5.847 $1.540$ (31)
Table 2: Relative precisions obtained in the parameter estimation of the sources listed in Table˜1 (where $\delta x$ denotes the relative precision in parameter $x$, defined as the $1\,\sigma$ width of the marginal posterior on $x$ normalised by the median value). The (s) superscripts indicates that masses have been converted in the source frame. For each sample, we recover the redshift from the luminosity distance $d_{\mathrm{L}}$ assuming the PLANCK18 cosmology. We also report $\Delta\Omega^{\mathrm{sky}}$ (Eq.˜16), which is $99\%$ of the sky-localisation area of each source in square degrees. In the last column, we provide the reference for the corresponding posterior distributions in Appendix˜F. Notably, $\delta m_{1}^{\mathrm{(s)}}$ and $\delta m_{2}^{\mathrm{(s)}}$ are a factor of $\sim 3$ orders of magnitude larger than the equivalent quantities computed in the detector frame, $\delta m_{i}^{\mathrm{(d)}}$. This is a consequence of the mass relation $m_{i}^{(d)}=(1+z_{\mathrm{r}})\,m_{i}^{(s)}$, which introduces the (much larger) uncertainty in $d_{\mathrm{L}}$ due to its relationship with $z_{\mathrm{r}}$ via our assumed cosmology.
@@ V.6.1 EMRIs
We investigate three EMRIs with mass ratios $\epsilon\in\{10^{-4},10^{-5},10^{-6}\}$ and parameters given by the fifth, first and second row of Table˜1 respectively. As confirmed by our horizon redshift study in Section˜V.4, these sources are indeed observable and each are unique in their own way. Source 1 has high spin and initial eccentricity $(a,e_{0})=(0.998,0.730)$, and represents a fairly typical EMRI, assuming that many MBHs in nature are spinning rapidly Babak:2017tow ; Barausse:2012fy . The more extreme mass-ratio of Source 2 causes its inspiral trajectory to evolve on a time-scale much slower than that of Source 1. This is reflected by its two-year inspiral starting deep in the strong field of the primary ($p_{0}\sim 2.12$) for $(a,e_{0})=(0.998,0.425)$, terminating $\sim 0.15$ away from the separatrix. Notice that the final eccentricity $e_{f}\sim 0.261$; the orbit is still quite eccentric close to plunge, and therefore lies in the upper tail of the eccentricity-at-plunge population for EMRIs formed via capture in Ref. Babak:2017tow . The final EMRI we consider, (Source 5) is a retrograde inspiral that (due to its mass ratio $\epsilon=10^{-4}$) begins in the weak-field regime ($p_{0}=26.19$) with high $e_{0}=0.8$ and moderate spin a=0.5a=-0.5, terminating near the separatrix with non-trivial eccentricity $e_{f}=0.195$. Based on a prior examination of EMRI astrophysics in Ref. Babak:2017tow , we believe these are characteristic asymmetric-mass binaries that are readily observable by LISA to plausible luminosity distances (and therefore redshifts).
@@
Measurement precisions on the parameters of these sources are shown in the corresponding rows of Table˜2. For Source 1, we see that we can constrain the intrinsic parameters to usual levels ($\Delta\theta/\theta\sim 10^{-6}$) for a source with an SNR of $30$. From the marginal posteriors (shown in Fig.˜28), we find that setting $\kappa=10^{-2}$ (retaining 235 modes) yields identical inference results to $\kappa=10^{-5}$ (1228 modes), with any systematic biases present being too small to identify. Similar results are obtained for Source 2 in terms of measurement precision (accounting for the lower SNR of this source), except for the improved determination of $e_{\mathrm{f}}$ which is expected due to the significant final eccentricity of this source. Correlations between parameters are more pronounced for this source than the other two, which is likely a feature of the mass ratio of this source leading to slower orbital parameter evolution. Similarly to Source 1, there is an absence of any systematic biases in Fig.˜29 between the posteriors for $\kappa=10^{-5}$ (809 modes) and $\kappa=10^{-2}$ (186 modes). For Source 3 5, the spin constraint are wider than for the other two sources, which is expected as the effects of spin are less pronounced for this source (as the inspiral spends a large fraction of the two-year duration in the weaker field). The same can not be said for eccentricity effects (which are still significant in the weak field), which is reflected in the recovery of $e_{\mathrm{f}}$. Once again, no systematic biases in Fig.˜31 are observable between posteriors for $\kappa=10^{-5}$ (928 modes) and $\kappa=10^{-2}$ (208 modes).
@@ V.6.2 IMRIs
As demonstrated in our horizon redshift study (Fig.˜17), IMRIs will be observable by LISA with high SNRs to high redshifts. To our knowledge, we have developed the first waveform model capable of modelling highly-eccentric IMRIs into rapidly-spinning MBHs (at the adiabatic order) that is sufficiently rapid for use in Bayesian inference studies. To demonstrate these capabilities and perform an initial investigation of IMRI posterior structure, we examine two IMRIs with mass ratios $\epsilon=10^{-2}$ and parameters given by row 3 and 4 in Table˜1. Due to the large mass ratio, we see that the orbital parameters evolve far more rapidly than for the EMRI sources we examined previously. Indeed, despite their high initial eccentricities $e_{0}=0.85$, the binaries have circularised significantly prior to plunge. Notably, the SNRs $\rho=\{500,200\}$ of these sources are also significantly larger, mainly due to the waveform amplitudes being proportional to $\mu\sim m_{2}$ (Eq.˜1). However, having such high SNRs may lead to more significant systematic biases when waveform models with incomplete mode content are used in parameter inference, as we explore below.
@@
For both Source 3 and Source 4, waveforms with $\kappa=10^{-5}$ (949 and 626 modes respectively) are sufficiently complete for performing parameter estimation with no observable biases, as shown in Figs.˜32 and 30. However, when attempting to infer the parameters with $\kappa=10^{-2}$ (not shown in Figures; 223 and 156 modes respectively) we observed statistically significant biases across all the IMRI parameters. Nearly all parameters were recovered with biases of $\gtrsim 3\,\sigma$ away from the truth (with $\sigma$ computed as the standard deviation of the posterior samples) that were readily identifiable compared to the scale of the posterior. Even for $\kappa=10^{-3}$ (412 and 291 modes respectively), we observe unacceptable biases in the luminosity distance and sky location on the order of $\sim 4\sigma$ and $\sim 7\sigma$ biases away from the true parameters respectively. Biases of this scale would propagate to population / cosmological analyses and may contaminate the results of these studies, especially when considering that luminosity distance biases will affect source-frame mass estimates as well. This behaviour is to be expected considering the higher SNRs of these sources — from Eq.˜15, we can expect mismatches of $10^{-3}$ (which $\kappa=10^{-3}$ roughly leads to) to be insufficient for the analysis of signals with SNRs of hundreds. The fact that extrinsic parameters are most affected by the omission of impactful waveform modes is also expected, as this has been previously identified in the analysis of comparable-mass case binaries observed with LISA Yi:2025pxe .
@@
We report measurement precisions of parameters of astrophysical interest for these sources in Table˜2. Despite the higher SNRs of these signals, their spins are recovered with larger uncertainties than for the prograde EMRI sources we examined. This can be understood in terms of the rapid evolution of the IMRI trajectories, which complete relatively few cycles in the strong-field regime where the effects of spin are most important. Eccentricity-at-plunge is also recovered more poorly; this has similar justification, in addition to the IMRI trajectories beginning at larger orbital separations and therefore entering the strong-field regime with lower eccentricities than the EMRI sources. Luminosity distances are recovered precisely (to better than $1\%$ precision) in both cases, which is to be expected as amplitude measurements depend strongly on the SNR. This has the interesting consequence that despite the recovery of detector-frame masses for the IMRI sources being less precise than for the EMRI sources, they still provide more precise recovery of source-frame masses due to the improvement in the luminosity distance measurement. Last, we note that the $99\%$ sky-localisation area for Source 3 ($\sim 17\,\mathrm{deg}^{2}$) is more than an order of magnitude larger than that of Source 4. This occurs because during early inspiral, the bulk of the mode spectrum for this source sits mainly at lower frequencies ($\lesssim 10^{-4}\,\mathrm{Hz}$) that are poorly measured by LISA. As the inspiral evolves during the observation, the mode spectrum shifts to higher frequencies (“entering” the lower end of the LISA band) and is measured well. While this still enables many parameters to be recovered precisely, the sky position of the source is determined by long-timescale variations in the LISA response function; as the waveform of this source only has significant SNR for part of the observation (such that less of these variations are measurable), this in turn leads to a poorer localisation of the source on the sky.
@@ VI.1 Summary of this work
By comparing with both the AAK and a quadrupolar variant of our waveform model, we demonstrated the importance of relativistic amplitudes and higher-order modes when performing SNR computations and parameter estimation. We found that SNRs computed with the AAK are more inaccurate for stronger-field high-mass systems, where the missing higher-order modes are most important and the semi-relativistic amplitudes of the AAK are least accurate. For all component masses considered, the AAK overestimated the SNR for large spins (by up to $60\%$ in the high-mass case), and exhibited larger SNR errors for larger eccentricities. This has ramifications for quantitative studies of the EMRI detection rate in the literature (such as Ref. Babak:2017tow ), which should be revisited with relativistic waveform models in order to obtain better rate estimates. Exploring inference biases under the linear signal approximation, we identify biases of $\sim 1\,\sigma$ with the AAK and $\sim 0.1\,\sigma$ with the quadrupolar relativistic model. While these biases may be identifiable in the parameter estimation of EMRIs, they are sufficiently small that these approximate models may be readily applicable in the search for EMRI signals in detector data (where any corresponding reductions in computational cost would be highly beneficial). An exploration of this possibility is left to future work.
To quantify and explore LISA’s sensitivity to GWs from asymmetric-mass binaries, we computed sky-averaged horizon (maximum) redshifts at which LISA can detect sources with varying component masses, spins and eccentricities. We found that LISA can detect EMRIs of mass ratio $10^{-6}$ ($10^{-4}$) at redshifts of up to 1 (5), and IMRIs of mass ratio $10^{-3}$ ($10^{-2}$) at redshifts of up to 10 (15), with only $\sim 10\%$ variations with respect to eccentricity. In the case of EMRIs with mass ratio $10^{-5}$, horizon redshifts varied from $0.5$ to $3$ as spin increased from 0 to $0.999$. For all configurations, we found that LISA was most sensitive to binaries with source-frame primary masses $m_{1}^{\mathrm{(s)}}\sim 10^{6}\,M_{\odot}$. Our horizon redshift results highlight that the detection and characterisation of asymmetric-mass binaries will enable LISA to effectively probe the evolution of MBH, intermediate-mass black hole and stellar-mass black hole populations over cosmic time, in line with LISA’s science objectives. The waveform model we have developed (as a component of the few framework) is unique in the respect that it is able to accurately reproduce the physical characteristics of this wide range of GW sources, providing the tools necessary for the exploration of LISA science with IMRI and EMRI observations.
@@ VI.2 Prospects for future work
Extension to generic orbits. The few framework is designed with generic, eccentric and precessing orbits in mind but some technical limitations must be overcome before such a model can be built. Of particular note is the increased difficulty in interpolating data for these inspirals; the parameter-space dimensionality increases to four which greatly expands the (already large) memory requirements for mode amplitude interpolation. Given that polar ($k\neq 0$) mode amplitudes must also be computed (increasing the total number of modes by more than an order of magnitude Fujita:2025pc ), extrapolating from the domain of our model leads to mode amplitude grids of $\sim 500\,\mathrm{GB}$ in size or more, which is several times too large to fit in the memory of existing GPU hardware. The overall cost of computing the underlying data for flux and amplitude interpolation also grows significantly as the higher dimensional parameter space necessitates sampling the space at more points. Furthermore, at each point in the parameter space the perturbation theory computation is more expensive as one must integrate over both an orbit’s radial and polar motions, and one must sum over both radial and polar frequency harmonics. This increases the computational cost of each mode which contributes to the flux by a factor of about 20203030, and increases the number of modes which contribute to the flux by a similar factor. The problem remains highly parallelizable, so will not be difficult to exploit efficient codes which can compute these data Drasco:2005kz ; Fujita:2009us ; vandeMeent:2017bcc ; Chen:2023lsa ; Kerachian:2023oiw . However, the sheer quantity of data needed means that care is needed in designing the grid spacing across the parameter space before committing significant computational resources.
Inspiral-merger-ringdown models. For typical EMRIs, very little SNR is accumulated during the merger and ringdown portion of the waveform. With recent work showing that GSF results can also be used to model IMRIs LeTiec:2011bk ; vandeMeent:2020xgc ; Wardell:2021fyy ; Warburton:2021kwk , it becomes more important to include the merger and ringdown Apte:2019txp ; Becker:2024xdi ; Lhost:2024jmw ; Leather:2025nhu . Previous approaches to computing these in perturbation theory involved computing the waveform using time-domain codes Taracchini:2014zpa ; Lim:2019xrb ; Strusberg:2025qfv and thus the results were not suitable for the few framework. To overcome this, recent work has cast the transition from inspiral to plunge and the plunging portions of the waveform within the multiscale framework Kuchler:2024esj ; Kuchler:2025hwx . Within this approach the online computation of the transition and the plunge involves solving new trajectory ODEs and the waveform is computed from a new set of amplitudes. Structurally, this is exactly the same as the present online inspiral computation and, as such, the results of these calculations can be readily incorporated into few.
@@
Post-adiabatic accuracy: including the second-order GSF. Current second-order GSF calculations are carried out via a multiscale framework Hinderer:2008dm ; Miller:2020bft ; pound2022black ; Mathews:2025nyb . Within this approach, the trajectory through the parameter space is computed using an extended version of Eq.˜2 that takes the form
@@
The post-adiabatic forcing functions $\hat{f}^{(1)}_{\alpha}$ are very computationally expensive to compute, but they can be pre-computed and interpolated in the same manner as the adiabatic forcing functions. Second-order GSF calculations also compute the second-order waveform amplitudes. It is expected that these are not necessary to include in EMRI waveforms Burke:2023lno , but they may be required when modelling low-$q$ IMRIs to avoid biases in parameter estimation. In that case, the second-order amplitudes will be interpolated along the sparse trajectory during online waveform generation alongside the adiabatic amplitudes. This means that online waveform generation of post-adiabatic waveforms will be almost as fast as adiabatic models. At present, the $\hat{f}^{(1)}_{\alpha}$ are only known for quasi-circular inspirals into a non-rotating black hole Wardell:2021fyy , and these results are ready to be incorporated into few (as was initially investigated in Ref. Burke:2023lno ). It is an active area of research to compute $\hat{f}^{(1)}_{\alpha}$ for more complex orbital configurations Spiers:2023cip ; Leather:2023dzj ; Osburn:2022bby ; Wei:2025lva . In addition to post-adiabatic contributions to the forcing functions, second-order corrections to the mode amplitudes may also be important for loud IMRIs (but not for EMRIs, see Ref. Burke:2023lno for discussion). Provided that GW fluxes are known to second order, the extension to second-order amplitudes is straightforward, being of the form
$\mathcal{A}_{\ell mkn}=\mathcal{A}_{\ell mkn}^{(0)}+\nu\mathcal{A}_{\ell mkn}^{(1)},$ (18)
@@
where $\mathcal{A}_{\ell mkn}^{(0)}$ are the adiabatic amplitudes described in Section˜III.2. Post-adiabatic amplitude corrections may also include contributions from the spin of the secondary object. Post-adiabatic accuracy: including the spin on the secondary. The effect of the spin on the secondary enters the waveform phase at post-adiabatic order. In terms of waveform generation, a spinning secondary introduces another forcing term to Eq.˜17 and the orbital frequencies are also modified. Both of these changes can readily be incorporated into few. At present, codes exist that can compute the necessary forcing terms for circular Harms:2015ixa ; Akcay:2019bvk ; Piovano:2020zin ; Mathews:2021rod ; Albertini:2024rrs , eccentric Skoupy:2021asz and generic orbits Skoupy:2023lih ; Drummond:2023wqc ; Piovano:2024yks . Last, while spin precession of the secondary object enters at second post-adiabatic order, it also contributes to mode phasing and amplitude evolution differently to other terms of this order Mathews:2025nyb ; the few framework provides the necessary tools for the importance of secondary spin precession to be examined in future work. With the exception of GW memory and generic orbits, work is already underway on all of these extensions (and we note that few currently can generate generic inspirals using a 55PN model Isoyama:2021jjd ). In addition to the above, few can also be extended to include environmental and beyond-GR physics. So long as any additional physics can be modelled as a forcing function added to Eq.˜2 or Eq.˜17 that do not depend on the orbital phase (or if such dependence can be transformed away VanDeMeent:2018cgn ; Lynch:2021ogr ; Lynch:2024hco ; Lynch:2024ohd ) it can readily be incorporated into our framework. As the few trajectory module now exists entirely in Python, if software for computing these modifications has already been developed it can be used directly to rapidly produce a new inspiral (and hence, waveform) model suitable for subsequent data analysis investigations. See, e.g., Refs. Copparoni:2025jhq and Speri:2024qak for first efforts in this direction.